OpenStudy (anonymous):

ABSOLUTE VALUE. Check for extraneous solutions. May someone please walk me through this? | 2x + 5 | = 3x + 4

OpenStudy (anonymous):

Bring all integers on one side of the equation. For example, on a problem like this: 5x - 4 = |2x + 1| Bring all integers on one side of the equation: 5x = 2x +1 +4 (add 4 to both sides) 5x = 2x + 5

OpenStudy (anonymous):

So on this particular question you are asking: | 2x + 5 | = 3x + 4 Bring all integers on one side of the equation: 2x + 5 - 5 = 3x + 4 - 5 (subtract 5 from both sides) 2x = 3x - 1

OpenStudy (anonymous):

Do you get that so far?

OpenStudy (anonymous):

@mangorox Question: When you were bringing all of the integers to one side... why is 5x on the left side of the equal sign and the other numbers on the right?

OpenStudy (anonymous):

@amoodarya What am I supposed to do with that information? o. o I'm confused...